Show that simple harmonic motion y(t) = C1 cos ωt + C2 sin ωt can be written as:

a. y(t) = A sin(ωt + ϕ0)

b. y(t) = A cos(ωt + ϕ1)


1
Expert's answer
2022-05-04T15:06:38-0400

At first, rewrite the function in the following way: "y(t)=\sqrt{C_1^2+C_2^2}(\frac{C_1}{\sqrt{C_1^2+C_2^2}}\cos\.wt+\frac{C_2}{\sqrt{C_1^2+C_2^2}}\sin\.wt)". Point out that "-1\leq\frac{C_1}{\sqrt{C_1^2+C_2^2}}\leq1" and "\left(\frac{C_1}{\sqrt{C_1^2+C_2^2}}\right)^2+\left(\frac{C_2}{\sqrt{C_1^2+C_2^2}}\right)^2=1" Therefore, there is such "\phi_0" that "\frac{C_1}{\sqrt{C_1^2+C_2^2}}=\cos\phi_0" and "\frac{C_2}{\sqrt{C_1^2+C_2^2}}=\sin\phi_0". It can be obtained from the properties of functions "\arccos", "\arcsin". We receive: "y(t)=(\cos\phi_0\,\cos\.wt+\sin\phi_0\sin\.wt)=\cos(wt-\phi_0)". Thus, we get: "y(t)=A\,\cos\.(wt-\phi_0)", where "A=\sqrt{C_1^2+C_2^2}" . After the change: "\phi_0\rightarrow-\phi_1" we receive formula b). After the change "w\rightarrow-w", "\phi_0\rightarrow\phi_0-\frac{\pi}2" in formula "y(t)=A\,\cos\.(wt-\phi_0)" we receive: "y(t)=A\,\cos\.(\frac{\pi}2-wt-\phi_0)=A\sin(wt+\phi_0)". It is formula a). .